OpenStudy (erinkb99):

Find the domain for the particular solution to the differential equation dy/dx = 3y/3x, with initial condition y(1) = 1.

OpenStudy (erinkb99):

I'm confused about the domain, but I found the separated the variables, took the antiderivatives, and found the particular solution

OpenStudy (erinkb99):

1/3y dy = 1/x dx (1/3)ln|x| = ln|x| +C

OpenStudy (erinkb99):

(1/3)ln|1| = ln|1|+C 0=0+c 0=C

jimthompson5910 (jim_thompson5910):

dy/dx = 3y/3x dy/3y = dx/3x taking the integral of both sides gives (1/3)*ln(|y|) = (1/3)*ln(|x|) + C Multiply everything by 3 to get (1/3)*ln(|y|) = (1/3)*ln(|x|) + C 3*(1/3)*ln(|y|) = 3*(1/3)*ln(|x|) + 3*C ln(|y|) = ln(|x|) + C then solve for y ln(|y|) = ln(|x|) + C |y| = e^(ln(|x|) + C) |y| = e^(ln(|x|)) * e^C |y| = e^C*e^(ln(|x|)) |y| = e^C*|x| Now we're given y(1) = 1 which means x = 1 and y = 1 |y| = e^C*|x| |1| = e^C*|1| 1 = e^C*1 e^C = 1 So the particular solution should be y = 1*|x| which simplifies to y = |x| For some reason, geogebra is saying the answer is y = x which is a bit odd. I'm not sure why it's not using absolute values

OpenStudy (erinkb99):

That means the domain is x>0?

jimthompson5910 (jim_thompson5910):

That could be a possibility

OpenStudy (erinkb99):

The answers are x > 0 x < 0 x ≠ 0 All real numbers

jimthompson5910 (jim_thompson5910):

I'm thinking all real numbers because either y = |x| like I showed above or y = x which is what geogebra got. Either way, the domain of each function is the set of all reals

OpenStudy (erinkb99):

Ok that makes sense, I was thinking more range

OpenStudy (erinkb99):

Thanks!

jimthompson5910 (jim_thompson5910):

glad to be of help